Glycolic acid, which is a monoprotic acid and a constituent in sugar cane, has a pKa of 3.9. A 25.0 mL solution of glycolic acid is titrated?
to the equivalence point with 35.8 mL of 0.020 M sodium hydroxide solution. What is the pH
of the resulting solution at the equivalence point?
1
Expert's answer
2018-03-19T08:46:58-0400
Solution:
This is a neutralization reaction:
NaOH + CH2(OH)COOH = CH2(OH)COONa + H2O
Execution of the ratio of which can find the pH value:
C1*V1 = C2*V2
-log[H+] = pH = 7.98.

Answer: 7.98